OpenStudy (anonymous):

Spelling mistakes in a text are either “nonword errors” or “word errors.” A nonword error produces a string of letters that is not a word, such as “the” typed as “teh.” Word errors produce the wrong word, such as “loose” typed as “lose.” Nonword errors make up 25% of all errors. A human proofreader will catch 80% of nonword errors and 50% of word errors. Reference: Ref 12-4 Of all the errors that the proofreader catches, what percent are word errors? A. 3.48% B. 37.5% C. 57.5% D. 65.2%

OpenStudy (anonymous):

lets use a good trick, make up some numbers

OpenStudy (anonymous):

i think i screwed that up. the wording is very confusing let me try again lets say there are 2000 errors, 25% are non word so 500 non word and the rest, 1500 word

OpenStudy (anonymous):

80% of 500 is 400 so 400 non word errors 50% of 1750 is 875 so 875 word errors are caught

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the total number of errors caught is 400 + 875 = 1275

OpenStudy (anonymous):

of those, 875 are word errors, so compute the percent \[\frac{875}{1275}\]

OpenStudy (anonymous):

i get 68% so either i made another mistake, or the answer is not in your list

OpenStudy (anonymous):

yeah i made a mistake 50% of 1500 is 750 and 750 + 400 = 1150

OpenStudy (anonymous):

\(\frac{750}{1150}=65.2\%\) rounded

OpenStudy (anonymous):

sorry i messed up before making up numbers when you are not given them is a good gimmick to use you can do it with decimals, but this way is usually easier